# FuseFlow: A Fusion-Centric Compilation Framework for Sparse Deep Learning on Streaming Dataflow

### v1

@@ 8.3. Fusion
We use FuseFlow to generate fused graphs for each model comparing the performance at different fusion levels against unfused configurations. Accelerating unoptimized code is rarely useful, making fusion an essential avenue for optimization. We show that it is important to identify the right fusion granularity to ensure meaningful performance gains. For SAE, GCN, and GraphSAGE, full fusion marks fusion regions around the end-to-end computations, while partial fusion focuses on fusion regions within encoder/decoder or individual layers. For GPT-3 with BigBird Attention, shape operators (e.g., reshape, transpose) naturally limit fusion, dividing decoders into multiple subsets. Full fusion across decoders combines these subsets across decoder boundaries.
As demonstrated in Figure 12, GPT-3 achieves up to ${\sim}2.7\times$ improvement with full fusion. Conversely, SAE, GCN, and GraphSAGE experience performance degradation under full fusion due to increased computational overhead from nested matrix multiplications. Partial fusion is more effective for these models, improving performance by up to ${\sim}2.6\times$ (GCN on OGB-collab) and ${\sim}3.9\times$ (GraphSAGE on OGB-mag). These results illustrate that optimal fusion granularity depends heavily on the model.
@@ 8.6. Dataflow Ordering
We evaluate the impact of dataflow ordering by varying the order of nested matrix multiplication (matmul)—a core operation in GCN and GraphSAGE—using KarateClub (Zachary, 1977). As shown in Figure 17, suboptimal orders cause up to ${\sim}29\times$ slowdown compared to the best. Leveraging the best order thus provides an end-to-end speedup of ${\sim}29\times$ for fused GCN and GraphSAGE models. Additionally, constraining each matmul kernel to the best dataflow order significantly reduces the design-space size by 68.5%-99.9%, as shown in Table 3. Without these constraints, GCN alone has an impractically large number of possible dataflows (estimated up to ${\sim}10^{15}$), so we limit the search space to $2\times 10^{8}$ configurations in FuseFlow.
@@ 9.1. Compiling Sparse Tensor Algebra to Dataflow
Several techniques have been proposed for compiling sparse tensor algebra to dataflow hardware. Closest to our work is the Custard compiler (Hsu et al., 2023). Custard compiles sparse tensor algebra expressions to SAM graphs with intra-layer iteration fusion (IIF). Moreover, the compiler for the Onyx chip (Koul et al., 2024) maps SAM graphs to physical sparse CGRA hardware. The SAMML dataflow graphs we target are an extension of SAM graphs with additional ML primitives. Unlike Custard, our work supports a different form of IIF through factored iteration and introduces cross-expression fusion (EKF).
An extension to the Spatial compiler (Rucker et al., 2021; Koeplinger et al., 2018) for Capstan hardware (Rucker et al., 2021; Prabhakar et al., 2017) compiles computations written in parallel patterns—a loop-based declarative language—to sparse dataflow hardware. Moreover, the Stardust (Hsu et al., 2025) compiler can compile high-level sparse tensor algebra languages to these parallel patterns. Stardust, like Custard, only supports tensor expressions and cannot compile entire sparse ML models or provide cross-expression fusion.
Finally, there is a class of work on compiling general-purpose C code to reconfigurable dataflow accelerators. Weng et al. (2022) describe a compiler from annotated sparse loops to their SPU hardware (Dadu et al., 2019), and Gobieski et al. (2023) presents a co-designed compiler from general code to a CGRA. Both works support sparse loops in their general-purpose input code but do not compile from higher-level sparse languages.
@@ 9.2. Fusion in Sparse Tensor Frameworks
The TACO compiler (Kjolstad et al., 2017) showed how to generate fused loops for sparse tensor algebra expressions on CPUs and GPUs (Senanayake et al., 2020). Later sparse tensor algebra compilers have expanded such fusion capabilities (Zhou et al., 2022) to expressions over additional data structures (Ahrens et al., 2023; Chou et al., 2018; Liu et al., 2024; Ye et al., 2022) and operations (Henry et al., 2021; Sundram et al., 2024; Root et al., 2024; Kovach et al., 2023). The fusion support in these compilers, however, is limited to tensor algebra expressions and CPU/GPU compilation (Yadav et al., 2022). Other specialized frameworks such as FusedMM (Rahman et al., 2021) and SeaStar (Wu et al., 2021) support sparse operator fusion but are limited to specific patterns (e.g., SDDMM+SpMM operations or GNN message-passing patterns). Zhou et al. (2022) introduce techniques that identify and avoid four common redundancy types in IIF for sparse tensor algebra. Their compiler, ReACT, is most similar to ours as it introduces a representation close to our fused Einsums to and generates factored iteration code. However, ReACT cannot fuse across multiple independent expressions. Building upon these works, we show how to fuse across independent Einsum expressions in an ML model and how to do so when compiling to dataflow machines.
The TeAAL framework (Nayak et al., 2023) presents a declarative language of cascaded Einsums to describe sparse tensor algebra accelerators and generates an accelerator simulator and performance model from that language. TeAAL represents multiple Einsum expressions similar to our work, but our work generates fused code across those Einsum expressions. While TeAAL is a tool for modeling dataflow accelerators, FuseFlow generates a program configuration or mapping to dataflow accelerators through the SAMML IR along with a simulation of that program.
@@ 9.3. Sparse ML Compilation and Frameworks
A few ML frameworks have been designed with support for sparse tensors and, hence, sparse ML models. Scorch (Yan et al., 2024) describes several techniques needed to implement a version of the PyTorch API that supports sparse as well as dense tensors. The MLIR Linalg + SparseTensor dialects (Bik et al., 2022) combined with the MLIR lowering from PyTorch to Linalg (Google, 2021) also provides a sparse ML framework for CPUs. Our compilation techniques complement these frameworks with a compilation path to sparse dataflow hardware. Domain-specific libraries like PyTorch Geometric (PyG) (Fey and Lenssen, 2019) and Deep Graph Library (DGL) (Wang et al., 2019) integrate sparse computation into specific applications, but they lack the generality needed for targeting a broader range of sparse models.
@@ Acknowledgments
Appendix A Appendix
A.1. Intra-expression Iteration Fusion Details
@@
A.2. Full Fusion Table for GraphSAGE Example
@@
A.3. Datasets
Table 4 shows a description of the datasets used to evaluate FuseFlow in Section 8.
Model Benchmark/Dataset MxN Sparsity
GCN/GraphSAGE Planetoid/Cora (Yang et al., 2016) 2708 x 1433 99.7%
GCN/GraphSAGE CitationFull/Cora_ML (Bojchevski and Günnemann, 2018) 2995 x 2879 99.8%
GCN/GraphSAGE CitationFull/DBLP (Bojchevski and Günnemann, 2018) 17716 x 1639 99.6%
GCN/GraphSAGE OGB-Collab (Hu et al., 2020) 235868 x 128 99.9%
GCN/GraphSAGE OGB-MAG (Hu et al., 2020) 1939743 x 128 99.9%
SAE ImageNet (Deng et al., 2009) 224 x 224 50%
SAE NIH-CXR (Wang et al., 2017) 1024 x 1024 50%
SAE LUNA16 (Setio et al., 2017) 512 x 512 50%
GPT-3 w/ BigBird IMDB (Bojchevski and Günnemann, 2018) 53.9%-86.5%*
Table 4. Datasets used for each model class with their corresponding sparsity levels. *Refers to attention mask sparsity.
A.4. Lowering Algorithm Implications
@@
Our proposed lowering method produces dataflow graphs with computations, along with their reductions, placed in their natural positions rather than deferring them to the end. In particular, The FuseFlow compiler generates factored iteration because it does not distribute multiplications across sums and does not construct a fully fused global iteration space. For our GraphSAGE example in Figure 9, we use background color shading to help visualize this placement and distinguish interleaved regions: blue shading highlights input iteration regions, while yellow shading highlights computation regions. Concretely, higher-order reducer primitives spatially appear earlier in the graph and generate coordinate streams that flow to stream joiners later in the graph. An abstracted version of this interleaving is shown in Figure 10 (right) with behavior equivalent to the factored fusion iteration space from Figure 5(b). On the other hand, Figure 10 (left) shows the generated SAM graphs from prior work (Hsu et al., 2023) with its behavior equivalent to the global iteration space in Figure 5(a). In this case, all computation is combined at the end, rather than interleaved. Figure 10 demonstrates how our sparse abstract machine dataflow graphs changed given the new lowering algorithm presented in this section. Concretely, for our GraphSAGE example, the SAM graph with global iteration space is shown in Figure 20(a), constrasting with the SAMML graph with factored iteration space as shown in Figure 20(b).
A.5. Full Cross-Expression Fusion Algorithm

### v2

@@ 8.3. Fusion
We use FuseFlow to generate fused graphs for each model comparing the performance at different fusion levels against unfused configurations. Accelerating unoptimized code is rarely useful, making fusion an essential avenue for optimization. We show that it is important to identify the right fusion granularity to ensure meaningful performance gains.
Fusion Configurations
We evaluate three fusion granularities: unfused (separate operations), partially fused, and fully fused. For graph models and SAE, partial fusion groups operations within each layer while full fusion merges all layers. For GPT-3, reshape operations act as fusion barriers; partial fusion groups operations between reshapes within decoder blocks, while full fusion additionally merges across decoder boundaries. See Appendix C in the Appendix for a visual breakdown of these exact fusion boundaries.
As demonstrated in Figure 12, GPT-3 achieves up to ${\sim}2.7\times$ improvement with full fusion. GCN and GraphSAGE experience performance degradation under full fusion due to increased computational overhead from nested matrix multiplications, so partial fusion remains more effective for these models (up to ${\sim}2.6\times$ for GCN on OGB-collab and ${\sim}3.9\times$ for GraphSAGE on OGB-mag). SAE achieves $1.94\times$ with full fusion but only ${\sim}1.01\times$ with partial fusion. Full fusion benefits from removing inter-layer materialization, but partial fusion offers limited benefit because each layer is dominated by a large sparse matrix multiplication, so fusing smaller subsequent operations provides minimal incremental gain. Models with similar patterns—large compute kernels followed by many smaller operations—will exhibit similar behavior.
@@
8.4. Comparison with Prior Dataflow Compilers
We compare FuseFlow against Custard (hsu2023sam) and Stardust (hsu2025stardust) (C+S), the two prior sparse dataflow compilers for general sparse tensor algebra. As discussed in Section 3, C+S only support IIF and require manual rewrites for cross-expression fusion. We evaluate on GCN with the OGB-Collab dataset (hu2020ogb) using identical simulator settings and hardware parameters across all configurations. The unfused baseline compiles each kernel independently, materializing all intermediate tensors to memory. The C+S (rewrite) configuration applies manual expression rewrites to C+S’s inputs that force fused output code within the constraints of C+S, only support for sparse tensor computations). FuseFlow uses our automatic cross-expression fusion techniques.
As shown in LABEL:tab:comp_speedup, C+S with handwritten rewrites achieves $1.97\times$ speedup over the unfused baseline. FuseFlow achieves an additional $1.33\times$ speedup over C+S, yielding $2.63\times$ total speedup. FuseFlow’s gains come from two sources: (1) automatic cross-expression fusion eliminates intermediate materializations that baseline C+S cannot fuse, and (2) factored iteration during code generation reduces coordinate processing overhead. Importantly, FuseFlow achieves this with less user effort since no manual expression rewrites as the input program to the compiler are required.
8.5. Sparsity Ablation Study
To isolate sparsity’s effect on fusion performance, we evaluate FuseFlow on 2-layer GCN using synthetic graphs (500 nodes, 128-dimensional features) with adjacency matrix sparsity varying from 50% to 95%. We test three graph structures (sparsity patterns): uniform random, power-law (scale-free networks), and block diagonal (clustered communities). Figure 15 shows that partial fusion achieves consistent speedups that increase with sparsity, as sparser matrices reduce coordinate processing overhead. Structured patterns (power-law, block diagonal) outperform uniform random due to better locality. In contrast, full fusion incurs slowdowns when coordination overhead dominates the reduced computation. These results confirm that optimal fusion granularity depends on both sparsity level and structure. FuseFlow scales with nonzero count rather than dense dimensions, which aligns with studies from prior sparse compilers (hsu2023sam; henry2021; kjolstad2017taco; kjolstad2020sparse).
Figure 15. Speedup over unfused baseline vs. adjacency matrix sparsity on a 2-layer GCN with synthetic graphs (500 nodes, 128 features) across three sparsity patterns: uniform random, power-law, and block diagonal.
@@ 8.8. Dataflow Ordering
We evaluate the impact of dataflow ordering by varying the order of nested matrix multiplication (matmul)—a core operation in GCN and GraphSAGE—using KarateClub (zachary1977information). As shown in Figure 18, suboptimal orders cause up to ${\sim}29\times$ slowdown compared to the best. Leveraging the best order thus provides an end-to-end speedup of ${\sim}29\times$ for fused GCN and GraphSAGE models. Additionally, constraining each matmul kernel to the best dataflow order significantly reduces the design-space size by 68.5%-99.9%, as shown in Table 4. Without these constraints, GCN alone has an impractically large number of possible dataflows (estimated up to ${\sim}10^{15}$), so we limit the search space to $2\times 10^{8}$ configurations in FuseFlow.
@@ 9.1. Compiling Sparse Tensor Algebra to Dataflow
Several techniques have been proposed for compiling sparse tensor algebra to dataflow hardware. Closest to our work is the Custard compiler (hsu2023sam). Custard compiles sparse tensor algebra expressions to SAM graphs with intra-layer iteration fusion (IIF). Moreover, the compiler for the Onyx chip (onyx) maps SAM graphs to physical sparse CGRA hardware. The SAMML dataflow graphs we target are an extension of SAM graphs with additional ML primitives. Unlike Custard, our work supports a different form of IIF through factored iteration and introduces cross-expression fusion (EKF).
An extension to the Spatial compiler (rucker2021capstan; spatial) for Capstan hardware (rucker2021capstan; plasticine) compiles computations written in parallel patterns—a loop-based declarative language—to sparse dataflow hardware. Moreover, the Stardust (hsu2025stardust) compiler can compile high-level sparse tensor algebra languages to these parallel patterns. Stardust, like Custard, only supports tensor expressions and cannot compile entire sparse ML models or provide cross-expression fusion.
Finally, there is a class of work on compiling general-purpose C code to reconfigurable dataflow accelerators. weng2022unifying describe a compiler from annotated sparse loops to their SPU hardware (dadu2019spu), and riptide presents a co-designed compiler from general code to a CGRA. Both works support sparse loops in their general-purpose input code but do not compile from higher-level sparse languages.
@@ 9.2. Fusion in Sparse Tensor Frameworks
The TACO compiler (kjolstad2017taco) showed how to generate fused loops for sparse tensor algebra expressions on CPUs and GPUs (senanayake2020). Later sparse tensor algebra compilers have expanded such fusion capabilities (zhou2022react) to expressions over additional data structures (ahrens2023looplets; chou2018; liu2024unisparse; sparsetir) and operations (henry2021; sundram2024recurrences; root2024shapes; kovach2023indexedstreams). The fusion support in these compilers, however, is limited to tensor algebra expressions and CPU/GPU compilation (yadav2022spdistal). Other frameworks such as FusedMM (rahman2021fusedmm) and SeaStar (wu2021seastar) support sparse operator fusion but are limited to specific patterns (e.g., SDDMM+SpMM operations or GNN message-passing patterns). zhou2022react introduce techniques that identify and avoid four common redundancy types in IIF for sparse tensor algebra. Their compiler, ReACT, is most similar to ours as it introduces a representation close to our fused Einsums to and generates factored iteration code. SparseLNR (dias2022sparselnr) extends TACO with selective fusion/distribution to balance complexity and locality. Both compilers generate factored iteration code similar to ours, but they cannot fuse multiple independent expressions or generate dataflow code. Building upon these works, we show how to fuse across independent Einsum expressions in an ML model and how to do so when compiling to dataflow machines.
The TeAAL framework (teaal) presents a declarative language of cascaded Einsums to describe sparse tensor algebra accelerators and generates an accelerator simulator and performance model from that language. TeAAL represents multiple Einsum expressions similar to our work, but our work generates fused code across those Einsum expressions. While TeAAL is a tool for modeling dataflow accelerators, FuseFlow generates a program configuration or mapping to dataflow accelerators through the SAMML IR along with a simulation of that program.
9.3. Relation to Classic Loop Optimizations
FuseFlow’s iteration-space transformations are the sparse dataflow analogues of classical loop optimizations studied extensively in prior compilers with sparse-loop optimizations (kjolstad2020sparse; bik1993; bik; venkat2015chillie) and the polyhedral compilation literature (feautrier1991dataflow; bondhugula2008practical; Strout2016SPF). Our intra-expression iteration fusion (IIF) corresponds to loop fusion, while dataflow order selection corresponds to loop interchange. However, unlike traditional polyhedral compilation that operates on dense affine iteration spaces, FuseFlow fuses sparse tensor algebra operations whose iteration spaces can be thought of as polyhedra with holes (kjolstad2017taco). In dataflow, we operate on compressed-coordinate streams with strict ordering constraints; these streams can be viewed as a linearization of the sparse iteration space (kovach2023indexedstreams; hsu2023sam). Our POG encodes constraints on the loop-ordering scheduling space, analogous to dependence polyhedra, while fusion tables—similar to certain fusion information contained in schedule trees—reshape streaming dataflow primitives into a fixed iteration policy that aligns with the POG rather than rearranging imperative loop nests.
@@ 9.4. Sparse ML Compilation and Frameworks
A few ML frameworks have been designed with support for sparse tensors and, hence, sparse ML models. Scorch (yan2024scorch) describes several techniques needed to implement a version of the PyTorch API that supports sparse as well as dense tensors. The MLIR Linalg + SparseTensor dialects (bik) combined with the MLIR lowering from PyTorch to Linalg (mpact) also provides a sparse ML framework for CPUs. Our compilation techniques complement these frameworks with a compilation path to sparse dataflow hardware. Domain-specific libraries like PyTorch Geometric (PyG) (pytorch-geometric) and Deep Graph Library (DGL) (dgl) integrate sparse computation into specific applications, but they lack the generality needed for targeting a broader range of sparse models.
@@ Acknowledgments
Appendix A Artifact Appendix
A.1. Artifact Abstract
This appendix describes how to set up and run the FuseFlow system, which includes programs compiled using the FuseFlow compiler and run on the Comal simulator. Our artifact provides a Docker image containing all required dependencies (Python, Rust, MLIR via LLVM, protobuf, etc.) and scripts to reproduce the experimental results reported in this paper. The artifact can be executed with any x86-64 machine with Docker, Python3, Git, and Bash support, at least 64 GB of RAM, and more than 200 GB of disk space.
A.2. Artifact Check-List (Meta-Information)
  • Data set: We use select datasets from the following sources(yang2016revisitingsemisupervisedlearninggraph; bojchevski2018deepgaussianembeddinggraphs; setio2017validation; wang2017chestxray; deng2009imagenet; hu2020ogb).
  • Run-time environment: Docker, Git, Python 3, and bash need to be installed on the local machine. We recommend proficiency in bash and git.
  • Hardware: Any conventional x86-64 CPU with at least 64 GB of RAM.
  • Metrics: Number of FLOPs, number of bytes transferred or accessed, search space size, latency in cycles and normalized.
  • Output: Terminal outputs, files, graphs (PDF figures).
  • How much disk space required (approximately)?: Approximately 200 GB of disk space would be sufficient.
  • How much time is needed to prepare workflow (approximately)?: About 5 human-minutes and 10-20 compute-minutes.
  • How much time is needed to complete experiments (approximately)?: The total time to complete all experiments is approximately 5 human-minutes and 96 compute-hours when measured on a Google Cloud C2-standard-60 instance (60 logical threads running on a Intel Xeon Gold 6253CL Processor with 240 GB memory).
  • Publicly available?: Yes, on Github at the fuseflow-artifact and on a publicly available archive Figshare DOI.
  • Code licenses (if publicly available)?: MIT License
  • Workflow framework used?: Docker
  • Archived (provide DOI)?: Yes, the reserved DOI is this Figshare DOI.
A.3. Description
A.3.1. How to Access
The code for this submission can be downloaded from the fuseflow-artifact repository. The repository includes a Dockerfile that can be used to build the Docker image for the full evaluation of the artifact. The artifact is also available at this reserved Figshare DOI.
A.3.2. Hardware Dependencies
We recommend using an x86-64 machine with at least 64 GB of RAM. The more RAM available, the less compute-hours each experiment will take. The Figure 12 benchmark script takes in as a parameter the number of workers that dictate how many simultaneous simulations to schedule. By default we use 2, but it can be scaled up with the available memory. Running with 3 workers, it peaked at 140 GB of memory. Our compute-time estimates are calculated on a machine with 240 GB of RAM.
A.3.3. Software Dependencies
The artifact requires a machine with Docker, Git, Python 3, and bash installed. We evaluated the artifact with the following configuration Debian 6.1, Docker 20.10.24+dfsg1, Python 3.11.2, and GNU bash 5.2.15(1)-release on an Intel-based machine.
A.3.4. Data sets
We use select datasets from the following sources(yang2016revisitingsemisupervisedlearninggraph; bojchevski2018deepgaussianembeddinggraphs; setio2017validation; wang2017chestxray; deng2009imagenet; hu2020ogb). The full set of datasets corresponding to each model can be found at Table 2.
A.4. Installation
To install, first clone the fuseflow-artifact repository to the local machine. Then build the Docker image with the following commands (the build can take up to 20 minutes):
### Clone via HTTPS ###
\$ git clone --recursive https://github.com/lrubens/fuseflow-artifact.git
\$ git submodule update --init --recursive
\$ docker build -t fuseflow-artifact .
The Docker container can be started with the following command within a bash terminal. This command will also print the container ID CONTAINER_ID.
\$ docker run -d -it --rm fuseflow-artifact bash
The container can be attached to by running:
\$ docker attach <CONTAINER_ID>
Once attached to the docker container, it is important not to not type exit in the docker terminal as this will kill the container. The proper way to exit the docker is the sequence CTRL-p, CTRL-q.
A.5. Experimental Workflow
The experimental workflow for this artifact includes running scripts in the Docker container to run experiments and generate figures in the paper. The detailed instructions can be found in the README.md within the repository.
A.6. Evaluation and Expected Results
Within the Docker container, run the following to generate all results:
### In Docker Container ###
\$ bash scripts/run_all_benchmarks.sh
# ctrl+p ctrl+q
Once the experiments finish, detach the container by pressing ctrl+p and ctrl+q. To copy the experiment results and figures from the container, move outside of the fuseflow-artifact repository on the local machine and run the following commands: The CONTAINER_ID is the same ID used to attach to the container. You may also retrieve the CONTAINER_ID again by running docker ps in your terminal. The results and figures will be copied to fuseflow-artifact/results.
### In the local machine ###
# Within fuseflow-artifact/
\$ bash scripts/extract_results.sh
The expected results in the fuseflow-artifact/results directory are:
fuseflow-artifact/results
|- figure12.pdf
|- figure13.pdf
|- figure14.pdf
|- figure16a.pdf
|- figure16b.pdf
|- figure17.pdf
|_ figure18.pdf
  • Figure 12: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure12.pdf and
    figure12_results.json. Verify that the results match the figure.
  • Figure 13: We do not provide artifact evaluation code to reproduce the results for the hardware validation as it requires access to proprietary Xilinx FPGA tools and takes too long to synthesize the Verilog hardware.
  • Figure 14: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure14.pdf and
    figure14_results.json. Verify that the results match the figure.
  • Figure 16: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure16a.pdf and
    figure16a_results.json. Verify that the results match the figure.
  • Figure 16: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure16b.pdf and
    figure16b_results.json. Verify that the results match the figure.
  • Figure 18: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure17.pdf and
    figure17_results.json. Verify that the results match the figure.
  • Figure 18: The reproduced figure and experimental results can be found in the fuseflow-artifact/results folder under figure18.pdf and
    figure18_results.json. Verify that the results match the figure.
Appendix B Appendix
B.1. Intra-expression Iteration Fusion Details
@@
B.2. Full Fusion Table for GraphSAGE Example
@@
B.3. Lowering Algorithm Implications
@@
Our proposed lowering method produces dataflow graphs with computations, along with their reductions, placed in their natural positions rather than deferring them to the end. In particular, The FuseFlow compiler generates factored iteration because it does not distribute multiplications across sums and does not construct a fully fused global iteration space. For our GraphSAGE example in Figure 10, we use background color shading to help visualize this placement and distinguish interleaved regions: blue shading highlights input iteration regions, while yellow shading highlights computation regions. Concretely, higher-order reducer primitives spatially appear earlier in the graph and generate coordinate streams that flow to stream joiners later in the graph. An abstracted version of this interleaving is shown in Figure 11 (right) with behavior equivalent to the factored fusion iteration space from Figure 5(b). On the other hand, Figure 11 (left) shows the generated SAM graphs from prior work (hsu2023sam) with its behavior equivalent to the global iteration space in Figure 5(a). In this case, all computation is combined at the end, rather than interleaved. Figure 11 demonstrates how our sparse abstract machine dataflow graphs changed given the new lowering algorithm presented in this section. Concretely, for our GraphSAGE example, the SAM graph with global iteration space is shown in Figure 21(a), constrasting with the SAMML graph with factored iteration space as shown in Figure 21(b).
B.4. Full Cross-Expression Fusion Algorithm